### v1

@@ 2 Notation and preliminaries
[…]
@@
An extended-real-value function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be proper if its domain $\mathrm{dom}(\varphi):=\{x\in\mathbb{R}^{n}:\varphi(x)<+\infty\}$ is nonempty. A proper function $\varphi$ is said to be closed if $\varphi$ is lower semi-continuous on $\mathbb{R}^{n}$. A function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be locally Lipschitz continuous around $x\in\mathrm{dom}(\varphi)$ with a modulus $L_{x}>0$ (resp., relative to $\mathrm{dom}(\varphi)$), if there exists $\delta_{x}>0$ such that $|\varphi(u)-\varphi(v)|_{2}\leq L_{x}\|u-v\|_{2}$ holds, for all $u,v\in\mathcal{B}(x,\delta_{x})$ (resp., for all $u,v\in\mathcal{B}(x,\delta_{x})\cap\mathrm{dom}(\varphi)$). A function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is said to be locally Lipschitz differentiable around $x\in\mathrm{dom}(\varphi)$ with a modulus $\tilde{L}_{x}>0$, if there exists $\delta_{x}>0$, such that $\varphi$ is continuously differentiable on $\mathcal{B}(x,\delta_{x})$ and $\|\nabla\varphi(u)-\nabla\varphi(v)\|_{2}\leq\tilde{L}_{x}\|u-v\|_{2}$ holds for all $u,v\in\mathcal{B}(x,\delta_{x})$.
@@ 2.1 Preliminaries on nonsmooth analysis
We first review some preliminaries on subdifferentials of nonconvex functions [15, 29, 34]. For a proper function $\varphi$, the Fréchet subdifferential at $x\in\mathrm{dom}(\varphi)$ is defined by
^$\widehat{\partial}\varphi(x):=\left\{y\in\mathbb{R}^{n}:\mathop{\lim\inf}\limits_{\begin{subarray}{c}z\to x\\ z\neq x\end{subarray}}\;\frac{\varphi(z)-\varphi(x)-\left\langle{y},{z-x}\right\rangle}{\|z-x\|_{2}}\geq 0\right\},$
@@
${\partial}\varphi(x):=\left\{y\in\mathbb{R}^{n}:\exists x^{k}\to x,~{}\varphi(x^{k})\to\varphi(x),~{}y^{k}\to y\text{ with $y^{k}\in\widehat{\partial}\varphi(x^{k})$ for each $k$}\right\}.$
@@
We set ^$\widehat{\partial}\varphi(x)=\partial\varphi(x)=\emptyset$ by convention when $x\notin\mathrm{dom}(\varphi)$, and define $\mathrm{dom}(\partial\varphi):=\{x:\partial\varphi(x)\neq\emptyset\}$. It is clear that ^$\widehat{\partial}(\alpha\varphi)(x)=\alpha\widehat{\partial}\varphi(x)$ holds for all $x\in\mathrm{dom}(\varphi)$ and $\alpha>0$, and ^$\widehat{\partial}\varphi(x)\subseteq{\partial}\varphi(x)$ holds for all $x\in\mathrm{dom}(\varphi)$. It holds that ^$\widehat{\partial}\varphi(x)=\{\nabla\varphi(x)\}$ if $\varphi$ is differentiable at $x$, and $\partial\varphi(x)=\{\nabla\varphi(x)\}$ holds when $\varphi$ is continuously differentiable around $x$ ([34, Exercise 8.8(b)]). For a convex $\varphi$, the Fréchet and limiting subdifferential both consist with the classical subdifferential of a convex function at each $x\in\mathrm{dom}(\varphi)$ ([34, Proposition 8.12]), that is,
@@
^$\widehat{\partial}\varphi(x)=\partial\varphi(x)=\{y\in\mathbb{R}^{n}:\varphi(z)-\varphi(x)-\left\langle{y},{z-x}\right\rangle\geq 0,\text{~{}for all~{}}z\in\mathbb{R}^{n}\}.$
@@
If $\varphi(x)=\varphi_{1}(x_{1})+\varphi_{2}(x_{2})$ for proper closed functions $\varphi_{1}:\mathbb{R}^{n_{1}}\to(-\infty,+\infty]$ and $\varphi_{2}:\mathbb{R}^{n_{2}}\to(-\infty,+\infty]$, where $x\in\mathbb{R}^{n}$ is expressed as $(x_{1},x_{2})$, then it holds that ^$\widehat{\partial}\varphi(x)=\widehat{\partial}\varphi_{1}(x_{1})\times\widehat{\partial}\varphi_{2}(x_{2})$ ([34, Proposition 10.5]). Moreover, for proper functions $\varphi_{1},~{}\varphi_{2}:\mathbb{R}^{n}\to(-\infty,+\infty]$, there holds ^$\widehat{\partial}\varphi_{1}(x)+\widehat{\partial}\varphi_{2}(x)\subseteq\widehat{\partial}(\varphi_{1}+\varphi_{2})(x)$ ([34, Corollary 10.9]), where the equality holds if $\varphi_{1}$ or $\varphi_{2}$ is differentiable at $x$ ([34, Section 8.8 Exercise]). For a function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$, the generalized Fermat’s Rule states that $0\in\widehat{\partial}\varphi(x^{\star})$ holds if $x^{\star}\in\mathrm{dom}(\varphi)$ is a local minimizer of $\varphi$ ([34, Theorem 10.1]).
@@
We finally review some properties of real-valued convex functions. Let $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ be convex. Then $\varphi$ is continuous on $\mathbb{R}^{n}$ [34, Theorem 2.35]. In view of this, one can verify that $\partial\varphi$ is outer semicontinuous, that is, $\bar{y}\in\partial\varphi(\bar{x})$ holds if there exists some $x^{k}\to\bar{x}$ and $y^{k}\to\bar{y}$ with $y^{k}\in\partial\varphi(x^{k})$ for each $k\in\mathbb{N}$. For any bounded $\mathcal{S}\subseteq\mathbb{R}^{n}$, the $\varphi$ is Lipschitz continuous on $\mathcal{S}$, and $\cup_{x\in\mathcal{S}}\partial\varphi(x)$ is nonempty and bounded ([6, Proposition 5.4.2]). The convex conjugate function $\varphi^{*}:\mathbb{R}^{n}\to[-\varphi(0),+\infty]$ is defined at $y\in\mathbb{R}^{n}$ as $\varphi^{*}(y):=\sup\{\left\langle{x},{y}\right\rangle-\varphi(x):x\in\mathbb{R}^{n}\}$. The conjugate $\varphi^{*}$ is a proper closed convex function [33, Theorem 12.2]. Furthermore, for any $x,y\in\mathbb{R}^{n}$, the following three statements are equivalent ([34, Proposition 11.3]): $\left\langle{x},{y}\right\rangle=\varphi(x)+\varphi^{*}(y)\Leftrightarrow y\in\partial\varphi(x)\Leftrightarrow x\in\partial\varphi^{*}(y)$. Therefore, it is clear that the Fenchel-Young Inequality, i.e., $\varphi(x)+\varphi^{*}(y)\geq\left\langle{x},{y}\right\rangle$, holds for all $x,y\in\mathbb{R}^{n}$, and it becomes an equality when $y\in\partial\varphi(x)$. Besides, from the Fenchel-Young Inequality, we know that $y\in\mathrm{dom}(\varphi^{*})$ if $y\in\partial\varphi(x)$ for some $x\in\mathbb{R}^{n}$.
2.2 KL property
@@
Definition 2.1 (KL property [1, 2]).
@@
A proper function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a KL function, if it satisfies the KL property at each point in $\mathrm{dom}(\partial\varphi)$. For connections between the KL property and the well-known error bound theory [27, 28, 31], we refer the interested readers to [11, 23]. The following proposition provides a framework for proving sequential convergence using the KL property.
Proposition 2.2.
[25, Proposition 2.7] Let $\Psi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to(-\infty,+\infty]$ be proper lower semicontinuous. Consider a bounded sequence $\{(u^{k},v^{k})\in\mathbb{R}^{n}\times\mathbb{R}^{m}:k\in\mathbb{N}\}$ satisfying the following three conditions:
@@
    (Continuity condition). The limit $\Psi_{\infty}:=\lim_{k\to\infty}\Psi(u^{k},v^{k})$ exists and $\Psi\equiv\Psi_{\infty}$ holds on $\Upsilon$, where $\Upsilon$ denotes the set of accumulation points of $\{(u^{k},v^{k}):k\in\mathbb{N}\}$.
@@ 2.3 Min-max problem
Problem (10) is called nonconvex-concave if for a fixed $x\in\mathcal{F}_{X}$, $\psi(x,\cdot)$ is concave, and for a fixed $y\in\mathcal{F}_{Y}$, $\psi(\cdot,y)$ is not convex. A point $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$ is said to be a saddle point of problem (10) if $\psi(x^{\star},y)\leq\psi(x^{\star},y^{\star})\leq\psi(x,y^{\star})$ holds for all $(x,y)\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$. The concept of saddle points is usually used to describe the optimality of the convex-concave min-max problems. Unfortunately, it is well-known that such points may not exist for nonconvex-concave min-max problems. Due to this, the notions of global and local min-max points are introduced in [21, Definition 9] and [21, Definition 14], respectively. Here we consider equivalent definitions of these two concepts given in [21, Remark 10 and Lemma 16] as follows.
Definition 2.3 (global min-max point).
[21, Remark 10] We say that $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$ is a global min-max point of problem (10), if $y^{\star}$ is a global maximizer of $\psi(x^{\star},\cdot)$ and $x^{\star}$ is a global minimizer of $\varphi$, i.e.,
(11) $\begin{cases}\varphi(x^{\star})\leq\varphi(x),&\text{for all $x\in\mathcal{F}_{X}$,}\\ \psi(x^{\star},y^{\star})\geq\psi(x^{\star},y),&\text{for all $y\in\mathcal{F}_{Y}$.}\end{cases}$
Definition 2.4 (local min-max point).
[21, Lemma 16] We say that $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$ is a local min-max point of problem (10), if $y^{\star}$ is a local maximizer of $\psi(x^{\star},\cdot)$ relative to $\mathcal{F}_{Y}$, and there exists an $\epsilon_{0}>0$ such that $x^{\star}$ is always a local minimizer of $\varphi_{\epsilon}$ for all $\epsilon\in(0,\epsilon_{0}]$ relative to $\mathcal{F}_{X}$, where the function $\varphi_{\epsilon}$ is defined as $\varphi_{\epsilon}(x):=\max\{\psi(x,y):y\in\mathbb{R}^{m},\|y-y^{\star}\|_{2}\leq\epsilon\}$.
In view of [21, Proposition 11], problem (10) always has a global min-max point if $\mathcal{F}_{X}$ is compact. A saddle point of problem (10) is also a local min-max point. For more discussions on their relationships, we refer the readers to [20, 21].

### v2

@@ 2 Notation and preliminaries
We begin with our preferred notations. We denote the Euclidean space of dimension $n$ and the set of nonnegative integers by $\mathbb{R}^{n}$ and $\mathbb{N}$. The $\ell_{1}$-norm, the $\ell_{2}$-norm and the inner-product of $\mathbb{R}^{n}$ are denoted by $\|\cdot\|_{1},~\|\cdot\|_{2}$ and $\left\langle{\cdot},{\cdot}\right\rangle$, respectively. We use $\mathcal{B}(x,\delta)$ to denote an open ball centered at $x\in\mathbb{R}^{n}$ with radius $\delta>0$, i.e., $\mathcal{B}(x,\delta):=\{z\in\mathbb{R}^{n}:\|x-z\|_{2}<\delta\}$. The Cartesian product of the sets $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ is denoted by $\mathcal{A}_{1}\times\mathcal{A}_{2}$. The distance from a vector $x\in\mathbb{R}^{n}$ to a set $\mathcal{A}\subseteq\mathbb{R}^{n}$ is denoted by $\mathrm{dist}(x,\mathcal{A}):=\inf\{\|x-y\|_{2}:y\in\mathcal{A}\}$, and we adopt $\mathrm{dist}(x,\emptyset)=+\infty$ for convention. The indicator function on a nonempty set $\mathcal{A}\subseteq\mathbb{R}^{n}$ is defined by
@@
An extended-real-value function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be proper if its domain $\mathrm{dom}(\varphi):=\{x\in\mathbb{R}^{n}:\varphi(x)<+\infty\}$ is nonempty. A proper function $\varphi$ is said to be closed if $\varphi$ is lower semi-continuous on $\mathbb{R}^{n}$. A function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be locally Lipschitz continuous around $x\in\mathrm{dom}(\varphi)$ with a modulus $L_{x}>0$ (resp., relative to $\mathrm{dom}(\varphi)$), if there exists $\delta_{x}>0$ such that $|\varphi(u)-\varphi(v)|_{2}\leq L_{x}\|u-v\|_{2}$ holds, for all $u,v\in\mathcal{B}(x,\delta_{x})$ (resp., for all $u,v\in\mathcal{B}(x,\delta_{x})\cap\mathrm{dom}(\varphi)$). A function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is said to be locally Lipschitz differentiable around $x\in\mathrm{dom}(\varphi)$ with a modulus $\tilde{L}_{x}>0$, if there exists $\delta_{x}>0$, such that $\varphi$ is continuously differentiable on $\mathcal{B}(x,\delta_{x})$ and $\|\nabla\varphi(u)-\nabla\varphi(v)\|_{2}\leq\tilde{L}_{x}\|u-v\|_{2}$ holds for all $u,v\in\mathcal{B}(x,\delta_{x})$. For a proper closed function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$, the proximal operator of $\varphi$ at $x\in\mathbb{R}^{n}$ is denoted by $\mathrm{prox}_{\varphi}(x)$ and defined as
$\mathrm{prox}_{\varphi}(x):=\arg\min\left\{\varphi(z)+\frac{1}{2}\|x-z\|^{2}_{2}:z\in\mathbb{R}^{n}\right\}.$ (11)
The proximal operator $\mathrm{prox}_{\varphi}$ takes an $x\in\mathbb{R}^{n}$ and maps it into a closed subset of $\mathbb{R}^{n}$, which may be empty, a singleton or a subset with multiple vectors. If for a given $x\in\mathbb{R}^{n}$, the function $z\to\varphi(z)+\frac{1}{2}\|x-z\|^{2}_{2}$ is coercive, then $\mathrm{prox}_{\varphi}(x)$ is nonempty (see, for example, (Beck-Amir:2017SIAM, , Theorem 6.4)). Moreover, if $\varphi$ is convex, then $\mathrm{prox}_{\varphi}(x)$ is a singleton for all $x\in\mathbb{R}^{n}$. In particular, for problem (1), the proximal operator $\mathrm{prox}_{f_{\mathcal{C}}}(x)$ is nonempty for all $x\in\mathbb{R}^{n}$ due to $f_{\mathcal{C}}(x)\geq 0$.
@@ 2.1 Preliminaries on nonsmooth analysis
We first review some preliminaries on subdifferentials of nonconvex functions PangJongShi-CuiYing:2021ModernNonconvexNondifferentiableOptimization ; Boris:2006Variational_analysis ; Rockafellar2004Variational . For a proper function $\varphi$, the Fréchet subdifferential at $x\in\mathrm{dom}(\varphi)$ is defined by
$\widehat{\partial}\varphi(x):=\left\{y\in\mathbb{R}^{n}:\mathop{\lim\inf}\limits_{\begin{subarray}{c}z\to x\\ z\neq x\end{subarray}}\;\frac{\varphi(z)-\varphi(x)-\left\langle{y},{z-x}\right\rangle}{\|z-x\|_{2}}\geq 0\right\},$
@@
${\partial}\varphi(x):=\{y\in\mathbb{R}^{n}:\exists x^{k}\to x,~\varphi(x^{k})\to\varphi(x),~\\ y^{k}\to y\text{ with $y^{k}\in\widehat{\partial}\varphi(x^{k})$ for each $k$}\}.$
@@
We set $\widehat{\partial}\varphi(x)=\partial\varphi(x)=\emptyset$ by convention when $x\notin\mathrm{dom}(\varphi)$, and define $\mathrm{dom}(\partial\varphi):=\{x:\partial\varphi(x)\neq\emptyset\}$. It is clear that $\widehat{\partial}(\alpha\varphi)(x)=\alpha\widehat{\partial}\varphi(x)$ for all $x\in\mathrm{dom}(\varphi)$ and $\alpha>0$. Moreover, $\widehat{\partial}\varphi(x)\subseteq{\partial}\varphi(x)$ for all $x\in\mathrm{dom}(\varphi)$. If $\varphi$ is differentiable at $x$, then $\widehat{\partial}\varphi(x)=\{\nabla\varphi(x)\}$, and if $\varphi$ is continuously differentiable around $x$, one also has $\partial\varphi(x)=\{\nabla\varphi(x)\}$ ((Rockafellar2004Variational, , Exercise 8.8(b))). For a convex $\varphi$, the Fréchet and limiting subdifferentials coincide with the classical subdifferential of a convex function at each $x\in\mathrm{dom}(\varphi)$ ((Rockafellar2004Variational, , Proposition 8.12)); namely,
@@
$\widehat{\partial}\varphi(x)=\partial\varphi(x)=\{y\in\mathbb{R}^{n}:\varphi(z)-\varphi(x)-\left\langle{y},{z-x}\right\rangle\geq 0,\text{~for all~}z\in\mathbb{R}^{n}\}.$
@@
If $\varphi(x)=\varphi_{1}(x_{1})+\varphi_{2}(x_{2})$ for proper closed functions $\varphi_{1}:\mathbb{R}^{n_{1}}\to(-\infty,+\infty]$ and $\varphi_{2}:\mathbb{R}^{n_{2}}\to(-\infty,+\infty]$, where $x\in\mathbb{R}^{n}$ is expressed as $(x_{1},x_{2})$, then it holds that $\widehat{\partial}\varphi(x)=\widehat{\partial}\varphi_{1}(x_{1})\times\widehat{\partial}\varphi_{2}(x_{2})$ ((Rockafellar2004Variational, , Proposition 10.5)). For proper functions $\varphi_{1},~\varphi_{2}:\mathbb{R}^{n}\to(-\infty,+\infty]$, there holds $\widehat{\partial}\varphi_{1}(x)+\widehat{\partial}\varphi_{2}(x)\subseteq\widehat{\partial}(\varphi_{1}+\varphi_{2})(x)$ ((Rockafellar2004Variational, , Corollary 10.9)), and equality occurs when $\varphi_{1}$ or $\varphi_{2}$ is differentiable at $x$ ((Rockafellar2004Variational, , Section 8.8 Exercise)). For a function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$, the generalized Fermat’s Rule states that $0\in\widehat{\partial}\varphi(x^{\star})$ if $x^{\star}\in\mathrm{dom}(\varphi)$ is a local minimizer of $\varphi$ ((Rockafellar2004Variational, , Theorem 10.1)).
@@
We finally review some properties of real-valued convex functions. Let $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ be convex. Then $\varphi$ is continuous on $\mathbb{R}^{n}$ (Rockafellar2004Variational, , Theorem 2.35). In view of this, one can verify that $\partial\varphi$ is outer semicontinuous, that is, $\bar{y}\in\partial\varphi(\bar{x})$ holds if there exists some $x^{k}\to\bar{x}$ and $y^{k}\to\bar{y}$ with $y^{k}\in\partial\varphi(x^{k})$ for each $k\in\mathbb{N}$. For any bounded $\mathcal{S}\subseteq\mathbb{R}^{n}$, the $\varphi$ is Lipschitz continuous on $\mathcal{S}$, and $\cup_{x\in\mathcal{S}}\partial\varphi(x)$ is nonempty and bounded ((Bertsekas:Convexoptimizationtheory, , Proposition 5.4.2)). The convex conjugate function $\varphi^{*}:\mathbb{R}^{n}\to[-\varphi(0),+\infty]$ is defined at $y\in\mathbb{R}^{n}$ as $\varphi^{*}(y):=\sup\{\left\langle{x},{y}\right\rangle-\varphi(x):x\in\mathbb{R}^{n}\}$. The conjugate $\varphi^{*}$ is a proper closed convex function (Rockafellar:70, , Theorem 12.2). Furthermore, for any $x,y\in\mathbb{R}^{n}$, the following three statements are equivalent ((Rockafellar2004Variational, , Proposition 11.3)): $\left\langle{x},{y}\right\rangle=\varphi(x)+\varphi^{*}(y)\Leftrightarrow y\in\partial\varphi(x)\Leftrightarrow x\in\partial\varphi^{*}(y)$. Therefore, it is clear that the Fenchel-Young Inequality, i.e., $\varphi(x)+\varphi^{*}(y)\geq\left\langle{x},{y}\right\rangle$, holds for all $x,y\in\mathbb{R}^{n}$, and it becomes an equality when $y\in\partial\varphi(x)$. In addition, from the Fenchel-Young Inequality, we know that $y\in\mathrm{dom}(\varphi^{*})$ if $y\in\partial\varphi(x)$ for some $x\in\mathbb{R}^{n}$.
2.2 KL property and semialgebra
@@
Definition 1(KL property Attouch-bolt-redont-soubeyran:2010 ; Attouch-Bolte:MP:2013 )
@@
A proper function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a KL function, if it satisfies the KL property at each point in $\mathrm{dom}(\partial\varphi)$. The following proposition provides a framework for full-sequential convergence analysis using the KL property.
Proposition 2.1
(Li-Shen-Zhang-Zhou:2022ACHA, , Proposition 2.7) (Sabach-Teboulle-Voldman:2018PAFA, , Theorem 4.3) Let $\Psi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to(-\infty,+\infty]$ be proper lower semicontinuous. A bounded sequence $\{(u^{k},v^{k})\in\mathbb{R}^{n}\times\mathbb{R}^{m}:k\in\mathbb{N}\}$ satisfies the following three conditions:
@@
    (Continuity condition). [^2] In (Sabach-Teboulle-Voldman:2018PAFA, , Theorem 4.3), this condition is replaced by $\lim\sup_{k\in\mathcal{K}\subseteq\mathbb{N}}\Psi(u^{k},v^{k})\leq\Psi(u^{\star},v^{\star})$, where $(u^{\star},v^{\star})$ is a limit point of a subsequence $\{(u^{k},v^{k})\}_{k\in\mathcal{K}}$. The limit $\Psi_{\infty}:=\lim_{k\to\infty}\Psi(u^{k},v^{k})$ exists and $\Psi\equiv\Psi_{\infty}$ holds on $\Upsilon$, where $\Upsilon$ denotes the set of accumulation points of $\{(u^{k},v^{k}):k\in\mathbb{N}\}$.
@@
Proper closed semialgebraic functions constitute a broad and important class of KL functions. A function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called semialgebraic if its graph $\mathrm{Graph}(\varphi):=\{(x,s)\in\mathbb{R}^{n}\times\mathbb{R}:s=\varphi(x)\}$ is a semialgebraic subset of $\mathbb{R}^{n+1}$, that is, there exist a finite many real polynomials $G_{ij}$, $H_{ij}:\mathbb{R}^{n+1}\to\mathbb{R}$ such that
$\operatorname{Graph}(\varphi)=\bigcup^{p}_{j=1}\,\bigcap^{q}_{i=1}\,\{z\in\mathbb{R}^{n+1}:G_{ij}(z)=0,~H_{ij}(z)<0\}.$
It is clear that any real polynomial functions is semialgebraic. More generally, the class of semialgebraic functions is stable under finite sums, products, and composition, and the convex conjugate of a semialgebraic function is semialgebraic. We refer the reader to Attouch-bolt-redont-soubeyran:2010 ; Bolte-Sabach-Teboulle:MP:2014 for more details on semialgebraic functions and the KL property.
@@ 2.3 Min-max problem
Problem (12) is called nonconvex-concave if for a fixed $x\in\mathcal{F}_{X}$, $\psi(x,\cdot)$ is concave, and for a fixed $y\in\mathcal{F}_{Y}$, $\psi(\cdot,y)$ is not convex. A point $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$ is said to be a saddle point of problem (12) if $\psi(x^{\star},y)\leq\psi(x^{\star},y^{\star})\leq\psi(x,y^{\star})$ holds for all $(x,y)\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$. The concept of saddle points is usually used to describe the optimality of the convex-concave min-max problems. Unfortunately, it is well-known that such points may not exist for nonconvex-concave min-max problems. Due to this, the notions of global and local min-max points are introduced in (Jin-Netrapalli-Jordan:2020ICML, , Definition 9) and (Jin-Netrapalli-Jordan:2020ICML, , Definition 14), respectively. Here we consider equivalent definitions of these two concepts given in (Jin-Netrapalli-Jordan:2020ICML, , Remark 10 and Lemma 16) as follows.
Definition 2(global min-max point)
(Jin-Netrapalli-Jordan:2020ICML, , Remark 10) Let $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$. We say that $(x^{\star},y^{\star})$ is a global min-max point of problem (12), if $y^{\star}$ is a global maximizer of $\psi(x^{\star},\cdot)$ and $x^{\star}$ is a global minimizer of $\varphi$, i.e.,
$\begin{cases}\varphi(x^{\star})\leq\varphi(x),&\text{for all $x\in\mathcal{F}_{X}$,}\\ \psi(x^{\star},y^{\star})\geq\psi(x^{\star},y),&\text{for all $y\in\mathcal{F}_{Y}$.}\end{cases}$ (13)
Definition 3(local min-max point)
(Jin-Netrapalli-Jordan:2020ICML, , Lemma 16) We say that $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$ is a local min-max point of problem (12), if $y^{\star}$ is a local maximizer of $\psi(x^{\star},\cdot)$ relative to $\mathcal{F}_{Y}$, and there exists an $\epsilon_{0}>0$ such that $x^{\star}$ is always a local minimizer of $\varphi_{\epsilon}$ for all $\epsilon\in(0,\epsilon_{0}]$ relative to $\mathcal{F}_{X}$, where the function $\varphi_{\epsilon}$ is defined as $\varphi_{\epsilon}(x):=\max\{\psi(x,y):y\in\mathbb{R}^{m},\|y-y^{\star}\|_{2}\leq\epsilon\}$.
Due to the lower semicontinuity of $\varphi$, problem (12) always has a global min-max point if $\mathcal{F}_{X}$ is compact. A saddle point of problem (12) is also a local min-max point. For more discussions on their relationships, we refer the readers to Jiang-Chen:2023SIAM-OPT ; Jin-Netrapalli-Jordan:2020ICML .